There are currently 50 million cars in a certain country, decreasing by 4.9% annually. How many years will it take for this country to have 27 million cars? Round to the nearest year.
Answer by stanbon(75887)

There are currently 50 million cars in a certain country, decreasing by 4.9% annually. How many years will it take for this country to have 27 million cars? Round to the nearest year.
----------------------
A(t) = Ao(1-0.049)^t
27 mil = 50 mil (0.951)^t
0.951^t = (27/50)
t[log0.951] = log[27/50]
t = 12.264 years
=================
Cheers,
Stan H.